given that 4,12,x,y,324,972 is a geometrical sequence. find the sum of x and y
Answer by Theo(13342)

geometris sequence formula is An = A1 * r^(n-1)

A1 is equal to 4, n = 1
A2 is equal to 12, n = 2

formula from A1 to A2 becomes 12 = 4 * r^(1) which becomes 12 = 4 * r

solve for r to get r = 12/4 = 3

A3 is therefore equal to 4 * 3^2 = 36
A4 is therefore equal to 4 * 3^3 = 108
A5 is therefore equal to 4 * 3^4 = 324
A6 is therefore equal to 4 * 3^5 = 972

your sequence is 4,12,36,108,324,972

x = 36
y = 108

x + y = 144

on a recursive basis, the formula is A[n+1] = A[n]*r

since r = 3, this means the next term in the sequence is the previous term multiplied by 3

start with 4 * 3 = 12 * 3 = 36 * 3 = 108 * 3 = 324 * 3 = 972

x is the third term in the sequence
y is the fourth term in the sequence.

x = 36
y = 108

sum of x and y = 144.